13. The arrivals and departures of planes at an airport is such that computer simulations is often used to model the Ideal
Answer #1

Answer:

Given,

n = 10

p1 = 1/3

p2 = 7/15

p3 = 1/5

Here it is a multinomial distribution

i.e.,

P(x1 = 3 , x2 = 6 , x3 = 1) = (10! / 3!*6!*1! ) * p1^3*p2^6*p3^1

substitute p1,p2,p3 values in above expression

= (10! / 3!*6!*1! ) * (1/3)^3 * (7/15)^6 * (1/5)^1

= 840*0.037*0.01033*0.2

= 0.0642